# Hierarchical Indexing for Retrieval-Augmented Opinion Summarization

### v1

@@ Abstract
We propose a method for unsupervised abstractive opinion summarization, that combines the attributability and scalability of extractive approaches with the coherence and fluency of Large Language Models (LLMs). Our method, HIRO, learns an index structure that maps sentences to a path through a semantically organized discrete hierarchy. At inference time, we populate the index and use it to identify and retrieve clusters of sentences containing popular opinions from input reviews. Then, we use a pretrained LLM to generate a readable summary that is grounded in these extracted evidential clusters. The modularity of our approach allows us to evaluate its efficacy at each stage. We show that HIRO learns an encoding space that is more semantically structured than prior work, and generates summaries that are more representative of the opinions in the input reviews. Human evaluation confirms that HIRO generates more coherent, detailed and accurate summaries that are significantly preferred by annotators compared to prior work.
@@ 1 Introduction
Early work on opinion summarization extracted reviewers’ sentiment about specific features Hu and Liu (2004) or identified salient sentences based on centrality Erkan and Radev (2004), while more recent methods have proposed extractive models that use learned feature spaces Angelidis et al. (2021); Basu Roy Chowdhury et al. (2022). Prior work on abstractive opinion summarization has almost exclusively either required costly supervision Bražinskas et al. (2021); Cattan et al. (2023) or has assumed that the number of input reviews is limited Coavoux et al. (2019); Bražinskas et al. (2020); Amplayo et al. (2021a, b); Iso et al. (2021). This defeats the point of a summary: a user could feasibly read 8 reviews in a reasonable period of time. A good summarization system should be scalable, since popular products online may receive thousands of reviews. It should also be attributable, offering some evidence to justify its output. Finally, it should generate summaries that are coherent and faithful to the input reviews.
Large Language Models (LLMs) have been shown to generate highly fluent summaries in the news domain Bhaskar et al. (2023) but are a flawed solution because current instruction-tuned models are not attributable, and because their context windows limit the number of reviews they are able to base their summaries on. Models with long context windows have been proposed Beltagy et al. (2020); Gu et al. (2022) but these are not currently instruction-tuned, and it has been shown that LLMs are biased toward information at the start and end of the input Liu et al. (2023).
Our approach, Hierarchical Indexing for Retrieval-Augmented Opinion Summarization (HIRO), identifies informative sentences using hierarchical indexing and then passes the selected sentences as input to a LLM, similar to retrieval-augmented generation (RAG, Lewis et al., 2020). This separation of concerns (selection and generation) allows us to combine the attributability and scalability of the discrete representation with the strong generative abilities of LLMs, leading both to a higher quality index and to more informative and coherent output summaries.
@@ 2 Related Work
Prior approaches to generating summaries of reviews broadly fall into two categories. Extractive methods generate summaries by selecting representative sentences from input reviews to use as the summary Erkan and Radev (2004); Angelidis et al. (2021); Basu Roy Chowdhury et al. (2022). These types of approach are scalable and inherently attributable, but result in summaries that are overly detailed and not coherent. Abstractive methods ‘read’ input reviews and generate novel language to use as the summary Bražinskas et al. (2020); Iso et al. (2021). The resulting summaries are more fluent and coherent, but most prior abstractive methods are only able to consider a limited number of input reviews, whether because of combinatoric complexity or a maximum input length.
@@
Prior work has investigated how to learn representations of data with richer structure and interpretability. Vendrov et al. (2016) propose learning an embedding space where the ordering of two pairs of samples could be inferred from their relative positions, but their method requires supervision of the correct ordering. Opper et al. (2023) describe a method for learning embeddings that explicitly include structural information, but they focus on representing structures, rather than on learning an ordering within the embedding space itself. Li et al. (2023) learn a tree-based index for passage retrieval concurrently with the dense embedding space, showing that this leads to improved retrieval performance. However, they focus on retrieving passages relevant to a query, rather than aggregating opinions.
@@
The idea of first selecting relevant parts of an input before generating output has been well studied (Kedzie et al., 2018; Puduppully et al., 2019; Amplayo et al., 2021b; Narayan et al., 2023, inter alia), and has been shown to be very effective when used in conjunction with LLMs in the form of retrieval-augmented generation (RAG, Lewis et al., 2020). Xu et al. (2023) found that retrieval augmentation is beneficial even when using models that can accept long inputs.
@@
Automatic evaluation of generated summaries is extremely challenging. Prior work has shown that ROUGE Lin (2004) scores correlate poorly with human assessments of summary quality Shen and Wan (2023); Clark et al. (2023); Aharoni et al. (2023). Some datasets are created automatically, with references that are not directly based on the input reviews Bražinskas et al. (2021). Modern summarization system outputs are now often preferred to human-written references Goyal et al. (2022); Bhaskar et al. (2023).
SummaC Laban et al. (2022) is a reference-free metric that uses an NLI model to evaluate the degree of support between a summary and the input documents, but it overly rewards trivial statements; using the obvious statement “The hotel was a building” as a summary for every entity achieves a near-perfect SummaC score of 99.9% on Space, a dataset of hotel reviews Angelidis et al. (2021).
@@ 4 Learning a Hierarchical Indexing
We base our encoder on Hercules Hosking et al. (2023b). Hercules uses an encoder-decoder architecture with a discrete hierarchical bottleneck to generate summaries. It is trained as a denoising autoencoder, and therefore needs to learn a representation that is both compressed enough to enable aggregation, but also expressive enough for the decoder to be able to generate meaningful output. These factors are in direct competition, with the compressed bottleneck leading to output that is generic and contains hallucinations. By contrast, HIRO uses an external LLM as the ‘decoder’, allowing us to focus on learning a representation that is useful for identifying informative opinions.
@@ 4.1 Method
The HIRO encoder module maps a single sentence $\mathbf{x}$ to a path $q_{1:D}$ through a discrete hierarchy, using a technique based on residual vector quantization Chen et al. (2010); Zeghidour et al. (2022); Hosking et al. (2023b).
First, we use a Transformer encoder followed by attention pooling Vaswani et al. (2017); Liu and Lapata (2019) to map a sequence of tokens $\mathbf{x}$ to a single dense embedding $\mathbf{z}\in\mathbb{R}^{\mathbb{D}}$. Then, we decompose $\mathbf{z}$ into a path through a latent discrete hierarchy $q_{1:D}$, where $q_{d}\in[1,K]$ are discrete ‘codes’ at each level $d$. Briefly, we induce a distribution over codes at each level $p(q_{d})$, parameterised by a softmax with scores $s_{d}$ given by the Euclidean distance from learned codebook embeddings to the residual error between the input and the cumulative embedding from all previous levels,
@@
where $\textbf{C}_{d}\in\mathbb{R}^{K\times\mathbb{D}}$ is a codebook which maps each discrete code to a continuous embedding $\textbf{C}_{d}(q_{d})\in\mathbb{R}^{\mathbb{D}}$. During training, we use the Gumbel reparameterization Jang et al. (2017); Maddison et al. (2017); Sønderby et al. (2017) to sample from the distribution $p(q_{d})$. During inference, we set $q_{d}=\operatorname*{arg\,max}s_{d}$.
@@
For each sentence in the training data, we construct positive pairs of semantically related sentences $\mathbf{x},\mathbf{x}_{+}$ as follows: given a random ‘query’ sentence $\mathbf{x}$ from the training data, we identify possible candidate ‘targets’ based on tf-idf similarity; then, we check whether each candidate is entailed by the query using an NLI model (DeBERTa v3, trained on Debiased NLI; He et al., 2021; Wu et al., 2022), and use the sentences which are labelled as entailed as positive targets $\mathbf{x}_{+}$. An example is shown in Figure 2. We do not use any ‘hard’ negatives; instead, during training we calculate the pairwise tf-idf similarity between all samples in a batch, and include only those samples with similarity below a threshold of $0.3$ as negatives $\mathbf{X}_{-}$. This prevents ‘false negatives’ being wrongly forced apart, and allows sentences that are topically related but not strictly equivalent to remain in the same high-level grouping within the index.
@@
We use a modified version of the InfoNCE training objective van den Oord et al. (2018) designed to bring the representations of a query $\mathbf{x}$ closer to those of a positive target $\mathbf{x}_{+}$, while pushing them apart from negative samples $\mathbf{X}_{-}$,
@@
where $\rho(\mathbf{x},\mathbf{x}_{+})$ is the tf-idf similarity between $\mathbf{x}$ and $\mathbf{x}_{+}$ that weights the confidence of the positive pairs, inspired by MarginMSE Hofstätter et al. (2021), and $\omega$ is a constant that controls the strength of the negative examples. The similarity function $s(\cdot,\cdot)$ is given by the mean dot product between the embedding of all subpaths $q_{1:d}$ for $d\leq D$,
@@ 4.2 Evaluation
We experiment using Space Angelidis et al. (2021), which consists of hotel reviews from TripAdvisor with 100 reviews per entity, as well as reference summaries created by annotators. We encode all the review sentences from the Space validation set, then calculate both the purity (the mean intra-cluster similarity) and colocation (the mean inter-cluster similarity) for the clusters of sentences assigned to each subpath $q_{1:d}$ for $d\leq 4$. Finally, we take the difference between the purity and colocation as an overall measure of the quality of the representation space.
We compare to Hercules Hosking et al. (2023b). We also compare to a baseline using recursive $k$-means over embeddings from a pretrained embeddings model (MiniLM, Reimers and Gurevych, 2019): we apply $k$-means to the embeddings, then calculate the residual errors between cluster centroids and embeddings, apply $k$-means to those errors, and repeat. All methods use the same number of clusters at each level ($k=12$).
Figure 3: Cluster quality by depth, as measured by the difference between cluster purity and colocation, according to NLI (solid line) and tf-idf (dashed line) similarity measures. HIRO learns a higher quality encoding space than comparison methods.
@@
We use a 6 layer Transformer, with token embeddings initialized from BERT base Devlin et al. (2019). [^1] We experimented with using BERT as the encoder but found no significant improvement, since the discrete encoding is the main bottleneck in the model. We set the codebook size $K=12$, with the number of levels $D=12$, based on development set performance. Other hyperparameters are given in Appendix B.
@@ 5.1 Method
Our content selection method identifies the parts of the hierarchy that are particularly popular for each entity, and extracts the corresponding clusters of sentences. This process is query-free; instead, we assign each subpath in the hierarchy $q_{1:d}$ a score based on its popularity within the indexed input reviews, and ‘retrieve’ all sentences that were mapped to the $k$ highest-scoring subpaths.
Our scoring function is inspired by tf-idf Jones (1972), which is designed to measure the importance of a particular term with respect to a set of baseline documents. We define the term popularity $\operatorname*{tp}(q_{1:d},e)$ of a path as the fraction of indexed reviews for entity $e$ which contain the subpath $q_{1:d}$,
@@ 5.2 Evaluation
We evaluate whether our method successfully selects clusters of sentences containing informative opinions, using reviews from Space, as in Section 4.2. First, we measure the overlap between retrieved clusters and oracle clusters constructed by selecting sentences with high overlap to the reference summaries, [^3] Space includes multiple reference summaries for each entity; we randomly select one when determining the oracle clusters. using the Adjusted Rand Index (ARI, Hubert and Arabie, 1985). Second, we evaluate the average prevalence Malon (2023) of sentences in retrieved clusters, which uses a NLI model (ALBERT, trained on VitC; Lan et al., 2020; Schuster et al., 2021) to evaluate how many input reviews support each retrieved sentence.
We compare HIRO to the clusters extracted by Hercules Hosking et al. (2023b), and the oracle upper bound. As a baseline, we apply $k$-means to MiniLM embeddings Reimers and Gurevych (2019), then extract the 25 sentences whose embeddings are closest to each centroid. For HIRO, we select the top $k=8$ subpaths for each entity, and set the smoothing constant $\alpha=6$.
@@ 6.1 Method
We generate extractive summaries by selecting the centroid of each cluster; we compute all pairwise ROUGE-2 scores between sentences in each cluster, and choose the sentence with highest average similarity to other cluster members. This approach is fully attributable, since each summary sentence is extracted verbatim from a review.
@@
We generate summaries one sentence at a time by passing the contents of a single cluster as input to an instruction-tuned LLM with a simple prompt that requests a single sentence that summarizes the main points. This leads to more fluent output that is still attributable, since each output sentence has an associated cluster of evidential sentences used to generate it.
@@
We generate summaries as a single document, by passing the sentences from all retrieved clusters for an entity to the LLM in one go. This gives summaries that are more coherent and less redundant, since the LLM has control over the whole summary. However, it is not possible to identify which cluster was used to generate each part of the summary, so the attributability is less granular.
The ideal balance between coherence and attributability is likely to vary by application.
@@
For the sentence-wise and document variants, we use Mistral 7B Instruct v0.2 to generate summaries from retrieved clusters. The LLM is queried in a zero-shot manner, and the prompts used are given in Appendix C. We sample with a temperature of $0.7$, and report the mean and standard deviation scores based on 3 samples.
6.2 Automatic Evaluation
$\texttt{SAP}=\texttt{prevalence}-\alpha~{}\texttt{genericness},$ (7)
@@ 6.2 Automatic Evaluation
Mistral 7B 5.3$\pm$0.1 19.6$\pm$0.4 41.3$\pm$1.3 34.3$\pm$3.3 24.2$\pm$0.8 1.9$\pm$0.0 12.6$\pm$0.0 17.3$\pm$0.2 17.6$\pm$0.4 8.5$\pm$0.2
Llama 2 7B 4.4$\pm$0.2 17.6$\pm$0.1 32.6$\pm$3.0 25.7$\pm$3.7 19.7$\pm$2.3 1.5$\pm$0.0 11.5$\pm$0.0 17.7$\pm$0.4 17.6$\pm$0.3 8.9$\pm$0.5
Llama 2 13B 6.5$\pm$0.5 17.6$\pm$0.4 22.9$\pm$1.7 21.9$\pm$1.8 12.0$\pm$0.8 2.9$\pm$0.1 12.9$\pm$0.0 10.7$\pm$0.1 13.9$\pm$0.1 3.8$\pm$0.2
HIROsent 4.5$\pm$0.1 18.2$\pm$0.0 36.2$\pm$0.8 20.1$\pm$0.5 26.2$\pm$0.9 3.5$\pm$0.0 14.1$\pm$0.1 14.6$\pm$0.3 6.9$\pm$0.1 11.2$\pm$0.3
HIROdoc 7.0$\pm$0.2 20.5$\pm$0.3 44.0$\pm$3.0 28.8$\pm$2.1 29.6$\pm$2.1 4.0$\pm$0.0 15.1$\pm$0.1 15.3$\pm$0.1 9.4$\pm$0.3 10.6$\pm$0.1
@@
Table 2: Results for automatic evaluation of summary generation. R-2 and R-L represent ROUGE-2/L F1 scores. Prev. refers to Prevalence, Gen. refers to Genericness, and SAP refers to Specificity-Adjusted Prevalence. We consider SAP to be our primary metric. The best scores for extractive and abstractive systems are shown in bold. Results for systems involving LLMs are based on 3 samples, with the mean and standard deviation shown. HIRO generates summaries with the best balance between prevalent opinions and specificity.

### v2

@@ Abstract
We propose a method for unsupervised abstractive opinion summarization, that combines the attributability and scalability of extractive approaches with the coherence and fluency of Large Language Models (LLMs). Our method, HIRO, learns an index structure that maps sentences to a path through a semantically organized discrete hierarchy. At inference time, we populate the index and use it to identify and retrieve clusters of sentences containing popular opinions from input reviews. Then, we use a pretrained LLM to generate a readable summary that is grounded in these extracted evidential clusters. The modularity of our approach allows us to evaluate its efficacy at each stage. We show that HIRO learns an encoding space that is more semantically structured than prior work, and generates summaries that are more representative of the opinions in the input reviews. Human evaluation confirms that HIRO generates significantly more coherent, detailed and accurate summaries.
@@ 1 Introduction
Early work on opinion summarization extracted reviewers’ sentiment about specific features (Hu and Liu, 2004) or identified salient sentences based on centrality (Erkan and Radev, 2004), while more recent methods have proposed extractive models that use learned feature spaces (Angelidis et al., 2021; Basu Roy Chowdhury et al., 2022). Prior work on abstractive opinion summarization has almost exclusively either required costly supervision (Bražinskas et al., 2021; Cattan et al., 2023) or has assumed that the number of input reviews is limited (Coavoux et al., 2019; Bražinskas et al., 2020; Amplayo et al., 2021a, b; Iso et al., 2021). This defeats the point of a summary: a user could feasibly read 8 reviews in a reasonable period of time. A good summarization system should be scalable, since popular products online may receive thousands of reviews. It should also be attributable, offering some evidence to justify its output. Paraphrasing Rashkin et al. (2023), we say that a statement $s$ is attributable to some evidence $E$, if a generic reader would agree that ‘According to $E$, $s$ is true’. Finally, it should generate summaries that are coherent and faithful to the input reviews.
Large Language Models (LLMs) have been shown to generate highly fluent summaries in the news domain (Bhaskar et al., 2023) but are a flawed solution because current instruction-tuned models are not attributable, and because their context windows limit the number of reviews they are able to base their summaries on. Models with long context windows have been proposed (Beltagy et al., 2020; Gu et al., 2022) but these are not currently instruction-tuned, and it has been shown that LLMs are biased toward information at the start and end of the input (Liu et al., 2023).
Our approach, Hierarchical Indexing for Retrieval-Augmented Opinion Summarization (HIRO), identifies informative sentences using hierarchical indexing and then passes the selected sentences as input to a LLM, similar to retrieval-augmented generation (RAG, Lewis et al., 2020). By separating the steps of content selection and generation, we can combine the attributability and scalability of the discrete representation with the strong generative abilities of LLMs, leading both to a higher quality index and to more informative and coherent output summaries.
@@ 2 Related Work
Prior approaches to generating summaries of reviews broadly fall into two categories. Extractive methods generate summaries by selecting representative sentences from input reviews to use as the summary (Erkan and Radev, 2004; Angelidis et al., 2021; Basu Roy Chowdhury et al., 2022). These types of approach are scalable and inherently attributable, but result in summaries that are overly detailed and not coherent. Abstractive methods ‘read’ input reviews and generate novel language to use as the summary (Bražinskas et al., 2020; Iso et al., 2021). The resulting summaries are more fluent and coherent, but most prior abstractive methods are only able to consider a limited number of input reviews, whether because of combinatoric complexity or a maximum input length.
@@
Other work has used specialised training data to train supervised models (e.g., Zhao et al., 2022; Cattan et al., 2023); however, such data is expensive to collect and is not generally available in every language, domain, or setting.
@@
Prior work has investigated how to learn representations of data with richer structure and interpretability. Vendrov et al. (2016) propose learning an embedding space where the ordering of two pairs of samples could be inferred from their relative positions, but their method requires supervision of the correct ordering. Opper et al. (2023) describe a method for learning embeddings that explicitly include structural information, but they focus on representing structures, rather than on learning an ordering within the embedding space itself. Li et al. (2023) learn a tree-based index for passage retrieval concurrently with the dense embedding space, showing that this leads to improved retrieval performance. However, they focus on retrieving passages relevant to a query, rather than aggregating opinions. Sarthi et al. (2024) use a pretrained embedding model to cluster related sentences then generate a summary using an LLM, and repeat this process iteratively to construct a tree-structured index over inputs. However, this process is expensive and must be repeated for each new set of inputs.
@@
The idea of first selecting relevant parts of an input before generating output has been well studied (Kedzie et al., 2018; Puduppully et al., 2019; Amplayo et al., 2021b; Narayan et al., 2023, inter alia), and has been shown to be very effective when used in conjunction with LLMs in the form of retrieval-augmented generation (RAG, Lewis et al., 2020). Xu et al. (2023) found that retrieval augmentation is beneficial even when using models that can accept long inputs. Wang et al. (2023) show that including an additional filtering or selection step to RAG is better than naively passing all retrieved documents as input.
@@
Automatic evaluation of generated summaries is extremely challenging. Prior work has shown that ROUGE (Lin, 2004) scores correlate poorly with human assessments of summary quality (Callison-Burch et al., 2006; Tay et al., 2019; Fabbri et al., 2021; Shen and Wan, 2023; Clark et al., 2023; Aharoni et al., 2023). Some datasets are created automatically, with references that are not directly based on the input reviews (Bražinskas et al., 2021). Modern summarization system outputs are now often preferred to human-written references (Goyal et al., 2022; Bhaskar et al., 2023).
SummaC (Laban et al., 2022) is a reference-free metric that uses an NLI model to evaluate the degree of support between a summary and the input documents, but it overly rewards trivial statements; using the obvious statement “The hotel was a building” as a summary for every entity achieves a near-perfect SummaC score of 99.9% on Space, a dataset of hotel reviews (Angelidis et al., 2021).
@@ 4 Learning a Hierarchical Indexing
We base our encoder on Hercules (Hosking et al., 2023b). Hercules uses an encoder-decoder architecture with a discrete hierarchical bottleneck to generate summaries. It is trained as a denoising autoencoder, and therefore needs to learn a representation that is both compressed enough to enable aggregation, but also expressive enough for the decoder to be able to generate meaningful output. These factors are in direct competition, with the compressed bottleneck leading to output that is generic and contains hallucinations. By contrast, HIRO uses an external LLM as the ‘decoder’, allowing us to focus on learning a representation that is useful for identifying informative opinions.
@@ 4.1 Method
The HIRO encoder module maps a single sentence $\mathbf{x}$ to a path $q_{1:D}$ through a discrete hierarchy, using a technique based on residual vector quantization (Chen et al., 2010; Zeghidour et al., 2022; Hosking et al., 2023b).
First, we use a Transformer encoder followed by attention pooling (Vaswani et al., 2017; Liu and Lapata, 2019) to map a sequence of tokens $\mathbf{x}$ to a single dense embedding $\mathbf{z}\in\mathbb{R}^{\mathbb{D}}$. Then, we decompose $\mathbf{z}$ into a path through a latent discrete hierarchy $q_{1:D}$, where $q_{d}\in[1,K]$ are discrete ‘codes’ at each level $d$. Briefly, we induce a distribution over codes at each level $p(q_{d})$, parameterised by a softmax with scores $s_{d}$ given by the Euclidean distance from learned codebook embeddings to the residual error between the input and the cumulative embedding from all previous levels,
@@
where $\textbf{C}_{d}\in\mathbb{R}^{K\times\mathbb{D}}$ is a codebook which maps each discrete code to a continuous embedding $\textbf{C}_{d}(q_{d})\in\mathbb{R}^{\mathbb{D}}$. During training, we use the Gumbel reparameterization (Jang et al., 2017; Maddison et al., 2017; Sønderby et al., 2017) to sample from the distribution $p(q_{d})$. During inference, we set $q_{d}=\operatorname*{arg\,max}s_{d}$.
@@
For each sentence in the training data, we construct positive pairs of semantically related sentences $\mathbf{x},\mathbf{x}_{+}$ as follows: given a random ‘query’ sentence $\mathbf{x}$ from the training data, we identify possible candidate ‘targets’ based on tf-idf similarity; then, we check whether each candidate is entailed by the query using an NLI model (DeBERTa v3, trained on Debiased NLI; He et al., 2021; Wu et al., 2022), and use the sentences which are labelled as entailed as positive targets $\mathbf{x}_{+}$. An example is shown in Figure 2. We do not use any ‘hard’ negatives; instead, during training we calculate the pairwise tf-idf similarity between all samples in a batch, and include only those samples with similarity below a threshold as negatives $\mathbf{X}_{-}$. We set this threshold to $0.3$ based on validation set performance. This prevents ‘false negatives’ being wrongly forced apart, and allows sentences that are topically related but not strictly equivalent to remain in the same high-level grouping within the index.
@@
We use a modified version of the InfoNCE training objective (van den Oord et al., 2018) designed to bring the representations of a query $\mathbf{x}$ closer to those of a positive target $\mathbf{x}_{+}$, while pushing them apart from negative samples $\mathbf{X}_{-}$,
@@
where $\rho(\mathbf{x},\mathbf{x}_{+})$ is the tf-idf similarity between $\mathbf{x}$ and $\mathbf{x}_{+}$ that weights the confidence of the positive pairs, inspired by MarginMSE (Hofstätter et al., 2021), and $\omega$ is a constant that controls the strength of the negative examples. The similarity function $s(\cdot,\cdot)$ is given by the mean dot product between the embedding of all subpaths $q_{1:d}$ for $d\leq D$,
@@ 4.2 Evaluation
We experiment using Space (Angelidis et al., 2021), which consists of hotel reviews from TripAdvisor with 100 reviews per entity, as well as reference summaries created by annotators. We encode all the review sentences from the Space test set, then calculate both the purity (the mean intra-cluster similarity) and colocation (the mean inter-cluster similarity) for the clusters of sentences assigned to each subpath $q_{1:d}$ for $d\leq 4$. Finally, we take the difference between the purity and colocation as an overall measure of the quality of the representation space.
We compare to Hercules (Hosking et al., 2023b) and a baseline using recursive $k$-means over embeddings from a pretrained embeddings model (MiniLM, Reimers and Gurevych, 2019). We apply $k$-means to the embeddings, then calculate the residual errors between cluster centroids and embeddings, apply $k$-means to those errors, and repeat. All methods use the same number of clusters at each level ($K=12$).
Figure 3: Cluster quality by depth, as measured by the difference between cluster purity and colocation for the Space test set, according to NLI (solid line) and tf-idf (dashed line) similarity measures. HIRO learns a higher quality encoding space than comparison methods.
@@
We use a 6 layer Transformer, with token embeddings initialized from BERT base (Devlin et al., 2019). [^1] We experimented with using BERT as the encoder but found no significant improvement, since the discrete encoding is the main bottleneck in the model. We set the codebook size $K=12$, with the number of levels $D=12$, based on validation set performance. Other hyperparameters are given in Appendix B.
@@ 5.1 Method
Our content selection method identifies the parts of the hierarchy that are particularly popular for each entity, and extracts the corresponding clusters of sentences. This process is query-free; instead, we assign each subpath in the hierarchy $q_{1:d}$ a score based on its popularity within the indexed input reviews, and ‘retrieve’ all sentences that were mapped to the $k$ highest-scoring subpaths. Our scoring function is inspired by tf-idf (Jones, 1972), which is designed to measure the importance of a particular term with respect to a set of baseline documents. We define the term popularity $\operatorname*{tp}(q_{1:d},e)$ of a path as the fraction of indexed reviews for entity $e$ which contain the subpath $q_{1:d}$,
@@ 5.2 Evaluation
We evaluate whether our method successfully selects clusters of sentences containing informative opinions, using reviews from Space, as in Section 4.2. First, we measure the overlap between retrieved clusters and oracle clusters constructed by selecting sentences with high overlap to the reference summaries, [^3] Space includes multiple reference summaries for each entity; we randomly select one when determining the oracle clusters. using the Adjusted Rand Index (ARI, Hubert and Arabie, 1985). Second, we evaluate the average prevalence (Malon, 2023) of sentences in retrieved clusters, which uses a NLI model (ALBERT, trained on VitC; Lan et al., 2020; Schuster et al., 2021) to evaluate how many input reviews support each retrieved sentence.
We compare HIRO to the clusters extracted by Hercules (Hosking et al., 2023b), and the oracle upper bound. As a baseline, we apply $k$-means to MiniLM embeddings (Reimers and Gurevych, 2019), then extract the 25 sentences whose embeddings are closest to each centroid. For HIRO, we select the top $k=8$ subpaths for each entity, and set the smoothing constant $\alpha=6$.
@@ 6.1 Method
We generate extractive summaries by selecting the centroid of each cluster; we compute all pairwise ROUGE-2 scores between sentences in each cluster, and choose the sentence with highest average similarity to other cluster members. This approach is inherently attributable, since each summary sentence is extracted verbatim from a review.
@@
We generate summaries one sentence at a time by passing the contents of a single cluster as input to an instruction-tuned LLM with a simple prompt that requests a single sentence that summarizes the main points. This leads to more fluent output that is likely to be attributable, since each output sentence has an associated cluster of evidential sentences used to generate it.
@@
We generate summaries as a single document, by passing the sentences from all retrieved clusters for an entity to the LLM in one go. This gives summaries that are more coherent and less redundant, since the LLM has control over the whole summary. However, it is not possible to identify which cluster was used to generate each part of the summary, and therefore more difficult to determine the attributability of the output.
The ideal balance between coherence and the granularity of associated evidence is likely to vary by application.
@@
For the variants that require an LLM, we use Mistral 7B Instruct v0.2 to generate summaries from retrieved clusters. Mistral 7B was chosen based on its qualitative performance during model development, but we compare using alternative models in Section 6.4. The LLM is queried in a zero-shot manner, and the prompts used are given in Appendix C. We sample with a temperature of $0.7$, and report the mean and standard deviation scores based on 3 samples.
@@
Zero-shot Mistral 7B 5.3$\pm$0.1 19.6$\pm$0.4 41.3$\pm$1.3 34.3$\pm$3.3 24.2$\pm$0.8 1.9$\pm$0.0 12.6$\pm$0.0 17.3$\pm$0.2 17.6$\pm$0.4 8.5$\pm$0.2
HIROsent+Mistral 7B 4.5$\pm$0.1 18.2$\pm$0.0 36.2$\pm$0.8 20.1$\pm$0.5 26.2$\pm$0.9 3.5$\pm$0.0 14.1$\pm$0.1 14.6$\pm$0.3 6.9$\pm$0.1 11.2$\pm$0.3
HIROdoc+Mistral 7B 7.0$\pm$0.2 20.5$\pm$0.3 44.0$\pm$3.0 28.8$\pm$2.1 29.6$\pm$2.1 4.0$\pm$0.0 15.1$\pm$0.1 15.3$\pm$0.1 9.4$\pm$0.3 10.6$\pm$0.1
@@
Table 2: Results for automatic evaluation of summary generation on the test splits. R-2 and R-L represent ROUGE-2/L F1 scores. Prev. refers to Prevalence, Gen. refers to Genericness, and SAP refers to Specificity-Adjusted Prevalence. ROUGE scores are no longer considered reliable (Callison-Burch et al., 2006; Tay et al., 2019; Fabbri et al., 2021), so we consider SAP to be our primary metric. The best scores for extractive and abstractive systems are shown in bold. Results for systems involving LLMs are based on 3 samples, with the mean and standard deviation shown. HIRO generates summaries with the best balance between prevalent opinions and specificity.
6.2 Automatic Evaluation
$\texttt{SAP}=\texttt{prevalence}-\alpha~{}\texttt{genericness},$ (7)